OpenStudy (logan13):

What is 4.62 expressed as the quotient of two integers in simplest form?

OpenStudy (logan13):

@MysticFury

OpenStudy (faiqraees):

x=4.62 100x =462 x = 462/100 x = 231/50

OpenStudy (welshfella):

well we can first write it as 462/100 can you reduce this to simplest form?

OpenStudy (logan13):

so divide 231/50

OpenStudy (welshfella):

no 231/50 is the simplest form - the answer

OpenStudy (logan13):

so wait

OpenStudy (logan13):

4.62

OpenStudy (welshfella):

there you go - that confirms that 231/50 is correct

OpenStudy (welshfella):

the GCF of 462 and 100 is 2 so dividing top and bottom by gave the 231/50

OpenStudy (logan13):

sooo

OpenStudy (welshfella):

you are not convinced?

OpenStudy (logan13):

so its 231/50

OpenStudy (logan13):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (logan13):

how would that make sense @welshfella